OpenStudy (anonymous):

how can i get the fraction out of this problem 4/3y+2=4y-4

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

have you tried a pair of pliers?

OpenStudy (amistre64):

to normalize a fraction; you multiply it by the denominator; everything by the denominator

OpenStudy (anonymous):

i can do everything but the fraction idk how to take it off

OpenStudy (amistre64):

(4/3)3y +2(3) = 4(3)y -4(3)

OpenStudy (amistre64):

4y +6 = 12y -12

OpenStudy (anonymous):

im going to try it can u plz wait a moment

OpenStudy (amistre64):

ill be here..

OpenStudy (anonymous):

i got 2.25 is that right

OpenStudy (amistre64):

lets see; 8y = 18 y = 18/8 = 9/4 = 2.25

OpenStudy (anonymous):

one answer is 5 the other is 6

OpenStudy (amistre64):

you check it by using it in the equation 4(9)/3(4)+2=4(9)/4-4 3+2 = 9-4 5=5 .. its good

OpenStudy (anonymous):

i multiply 4*2.5 i get 10 and then i subtract 4 and its 6

OpenStudy (amistre64):

yeah, try using the number you foound tho....2.25

OpenStudy (anonymous):

ooo thats wat i did wrong. thank u very much i really am thankful. i got a question about this site

OpenStudy (anonymous):

ooo thats wat i did wrong. thank u very much i really am thankful. i got a question about this site

OpenStudy (amistre64):

:) whats the question?

OpenStudy (anonymous):

is this site really free

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

does anyone ask you for a payment

OpenStudy (amistre64):

depends on your definition of free; it costs me a few hours of sleep ;)

OpenStudy (amistre64):

there is no charge for any services at this site

OpenStudy (anonymous):

im really shocked that people really help students out with their hw with out charging

OpenStudy (anonymous):

yah

OpenStudy (amistre64):

well, its either that or sit around and do nothing ....

OpenStudy (amistre64):

i use it to keep myself 'math aware'

OpenStudy (anonymous):

this is going to be my new favorite site.thanks to u =D

OpenStudy (amistre64):

well, good luck with it then ;)

OpenStudy (anonymous):

thank u have a good night